Circular No. 5778

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

VY AQUARII J. A. Mattei, AAVSO, reports that this cataclysmic variable is having

its first known outburst since 1990 June 30 (IAUC 5046), when it reached m_{v} = 10.5. Visual magnitude estimates: Apr. 24.59 UT, [13.5 (W. Albrecht, Pahala, HI); 24.70, [13.0 (A. F. Jones, Nelson, New Zealand); 25.57, 10.0 (Albrecht); 28.73, 10.3 (Jones); 29.68, 10.3 (Jones).

1993 HO_{1}

R. H. McNaught, Siding Spring Observatory, reports his discovery, on a U.K. Schmidt Telescope plate taken by K. S. Russell, of an asteroidal object with unusual motion. Follow-up observations were made on Apr. 26 by McNaught with the Uppsala Southern Schmidt and on Apr. 29 by C. P. Cass with the UKST. The following measurements are by McNaught:

1993 UT  α_{2000}  δ_{2000}  V
Apr. 20.67001  14^{h}55^{m}17.^{s}44  −20^{o}34′17′′.6  17.5
20.70473  14 55 13.93  −20 34 18.8
26.65900  14 45 30.41  −20 32 59.2  17
26.71528  14 45 24.67  −20 32 55.6
29.72234  14 40 35.62  −20 29 39.0

Orbital elements by G. V. Williams, Center for Astrophysics:

T = 1992 Dec. 18.867 TT  ω = 104.^{o}965
e = 0.41781  Ω = 22.942 }2000.0
q = 1.15895 AU  i = 5.926

a = 1.99068 AU n^{o} = 0.350914 P = 2.81 years

1993TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Apr. 23 14^{h}51.^{m}49 −20^{o}34.^{′}7 0.655 1.649 165.^{o}8 8.^{o}6 17.3
28 14 43.35 −20 31.8 0.674 1.677 171.8  4.9 17.2
May 3 14 35.54 −20 24.6 0.698 1.705 175.1  2.9 17.2
8 14 28.38 −20 14.6 0.728 1.733 171.8  4.8 17.4
13 14 22.08 −20 03.1 0.762 1.760 166.3  7.8 17.7
18 14 16.78 −19 51.5 0.802 1.788 160.5 10.9 18.0
23 14 12.59 −19 40.9 0.847 1.815 154.9 13.7 18.2
28 14 09.54 −19 32.3 0.896 1.843 149.5 16.2 18.5
June 2 14 07.59 −19 26.4 0.950 1.870 144.3 18.4 18.7

1993 April 30 Daniel W. E. Green